OpenStudy (anonymous):

Find the fifth roots of 243(cos 260° + i sin 260°)

OpenStudy (tkhunny):

Have you considered DeMoivre?

OpenStudy (anonymous):

Idk how to do that :l @tkhunny

OpenStudy (anonymous):

De Moivre's Theorem states that (cos x + i sin x)^n=cos nx + i sin nx, if I'm not mistaken.

OpenStudy (tkhunny):

\([243(cos 260° + i sin 260°)]^{1/5} =\) \(243^{1/5}\left[cos\left(\dfrac{x + 2k\pi}{5}\right) + i\sin\left(\dfrac{x + 2k\pi}{5}\right)\right]\) for k = 0, 1, 2, 3, 4 Really, just fill in the values for k.

OpenStudy (anonymous):

3[cos(x+2*0pi/5) + isin(x+2*0pi/5)]

OpenStudy (tkhunny):

Sorry, \(x = 260º\). I forgot to fill that in.

OpenStudy (anonymous):

What goes on now?

OpenStudy (anonymous):

3[cos(260+2) + isin(260+2)] 3[cos(262) + isin(262)]

OpenStudy (anonymous):

oh no! it will be 3[cos(260/5) + isin(260/5)

OpenStudy (tkhunny):

Please remember your Order of Operations. x + 2pi/5 is absolutely NOT the same as (x+2pi)/5 For k = 0, you have (260+0)/5 For k = 1, you have (260+360)/5 and so on. Normally, I'd prefer to use \(2\pi\) rather than 360º, but that 260º is a bit annoying. You could convert that to \(\dfrac{13\pi}{9}\) if you would like!

OpenStudy (anonymous):

3[(cos52) + isin(52)]

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

so when k = 2, it is 260 + 720?

OpenStudy (tkhunny):

Yup. That's the first one. Four more to go.

OpenStudy (tkhunny):

You are beginning to get the idea.

OpenStudy (anonymous):

So far: 3[cos(52) + isin(52)] 3[cos(124) + isin(124)] @tkhunny

OpenStudy (tkhunny):

Each new value should be different by 360/5 degrees. 360/5 = 72 124-52 = 72 I believe you have it!

OpenStudy (anonymous):

I've got it! 3[cos(52) + isin(52)] 3[cos(124) + isin(124)] 3[cos(196) + isin(196)] 3[cos(268) + isin(268)] 3[cos(340) + isin(340)] @tkhunny

OpenStudy (anonymous):

I wish I could give you best answer 10 times over! @tkhunny

OpenStudy (tkhunny):

Excellent work!